Calculate the break even point in dollars and units for a product that has:

Expected Unit Sales at launch:10,000 units

Selling Price/Unit: $9.99 Fixed Costs: $50,000

Variable Cost/Unit: $2.50
Answer #1

Contribution margin=Sales-Variable cost

=(9.99-2.5)=$7.49 per unit

Breakeven=Fixed expenses/Contribution margin

=50,000/7.49

=6675.57 units(Approx)

=(6675.57*9.99)

=$66688.92(Approx).